He spends 18 hours to finish 40% of a job, how lond will it take to finish the whole job?
Answer by mananth(16949)

40 % of x = 18
40/100 *x=18
multiply by 100
40x= 1800
/40
x= 1800/40
=45 hours
..
m.ananth@hotmail.ca